In a basketball game, Marlene made 16 field goals. Each of the field goals were worth either 2 points or 3 points, and Marlene scored a total of 39 points from the field goal.
Part A: Let x represent the number of two-point field goals and y represent the number of three-point field goals. Write a system of equations in terms of x and y to model the situation.
Part B: How many three-point field goals did Marlene make in the game?

Answer by nerdybill(7384)

In a basketball game, Marlene made 16 field goals. Each of the field goals were worth either 2 points or 3 points, and Marlene scored a total of 39 points from the field goal.
Part A: Let x represent the number of two-point field goals and y represent the number of three-point field goals. Write a system of equations in terms of x and y to model the situation.
.
x+y = 16 (equation 1)
2x+3y = 39 (equation 2)
.
Part B: How many three-point field goals did Marlene make in the game?
.
solve equation 1 for x:
x+y = 16
x = 16-y
substitute above into equation 2:
2x+3y = 39
2(16-y)+3y = 39
32-2y+3y = 39
32+y = 39
y = 7 (number of three-point goals)